The product of two consecutive integers is 2 less than the square of the lesser integer. Find the greater of the two integers
Answer by TimothyLamb(4379)

x = smaller integer
y = larger integer
---
y = x + 1
xy = xx - 2
---
xy = xx - 2
x(x + 1) = xx - 2
xx + x = xx - 2
x = -2
y = -1
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---